Six years ago Olwethu lent Happy R150 000 on condition that he would pay her back in nine years time. The applicable interest rate is 15,5% per year, compounded monthly. Happy also owes Olwethu another amount of R250 000 that he has to pay back six years from now for a loan that earned interest at 16,4% per year, compounded semi-annually. Happy asks Olwethu if he an settle both his debts three years from now. The total amount that Happy will have to pay Olwethu three years from now is

1
Expert's answer
2022-03-16T15:53:39-0400

"A = P(1 + \frac{r}{n})^{nt}\\=150000(1+\frac{0.155}{12})^{9×12}=599863.90,\\

250000(1+\frac{0.164}{6})^{3×6}=401147.02,\\


599863.90+401147.02=1001010.92"